# Training with Pseudo-Code for Instruction Following

### v1

@@ 3.3.1 Instruction-Following
[…]
And Chain Selection Selection and Chain
1 2 1 2 $\geq$3 2 $\geq$3
Llama 3.1 8B-NL 0.61 0.37 0.30 0.44 0.33 0.25 0.29 0.24
Llama 3.1 8B-PC 0.63 0.39 0.37 0.40 0.35 0.31 0.44 0.31
Mistral 7B v0.3-NL 0.51 0.29 0.25 0.35 0.25 0.23 0.25 0.20
Mistral 7B v0.3-PC 0.56 0.35 0.32 0.39 0.31 0.27 0.24 0.22
Qwen2.5 7B-NL 0.61 0.40 0.37 0.47 0.37 0.24 0.37 0.24
Qwen2.5 7B-PC 0.64 0.48 0.41 0.47 0.45 0.38 0.42 0.31
Granite 8B Code-NL 0.53 0.30 0.28 0.44 0.29 0.18 0.19 0.22
Granite 8B Code-PC 0.58 0.36 0.34 0.38 0.32 0.26 0.27 0.19
Granite 3.1 8B-NL 0.62 0.41 0.36 0.46 0.35 0.24 0.39 0.26
Granite 3.1 8B-PC 0.65 0.44 0.42 0.47 0.42 0.41 0.33 0.30
Chained and Nested Instructions: We use ComplexBench  (Wen et al., 2024) to study how models perform in the presence of conjunctions (And), sequential instructions that are simple chains (Chain), conditional sequential operations (Sequential) at varying depths of nesting and, combinations of such instructions (eg. selections and chain). As can be seen in Table 4, models trained with pseudo-code instructions consistently perform better than those trained with natural language instructions. This holds true for almost all instruction-categories as well as nesting-depths. It is also interesting to note that doing well on a shallow nesting-depth does not automatically mean that the performance trends follow on deeper nesting-depths (eg: LLama 3.1 8B and Granite 8B Code NL models on Selection tasks at nesting levels $11$ vs the rest).
Model JSON Format
Markdown
Format
Bullet
Format
Punctuation Length Start with End with Template
Llama 3.1 8B-NL 0.13 0.66 0.45 0.43 0.27 0.17 0.34 0.39
Llama 3.1 8B-PC 0.18 0.81 0.57 0.49 0.26 0.18 0.27 0.40
Mistral 7B v0.3-NL 0.11 0.58 0.45 0.23 0.24 0.15 0.25 0.23
Mistral 7B v0.3-PC 0.16 0.68 0.52 0.29 0.27 0.19 0.27 0.34
Qwen2.5 7B-NL 0.18 0.66 0.49 0.38 0.30 0.15 0.25 0.48
Qwen2.5 7B-PC 0.18 0.74 0.58 0.40 0.28 0.19 0.28 0.47
Granite 8B Code-NL 0.13 0.68 0.42 0.35 0.28 0.14 0.30 0.28
Granite 8B Code-PC 0.18 0.70 0.53 0.42 0.29 0.21 0.31 0.33
Granite 3.1 8B-NL 0.18 0.64 0.53 0.32 0.29 0.18 0.34 0.40
Granite 3.1 8B-PC 0.22 0.77 0.59 0.46 0.27 0.18 0.28 0.48
[…]
Hard Satisfaction Rate (HSR) % Soft Satisfaction Rate (SSR) %
Model L1 L2 L3 L4 L5 Avg. L1 L2 L3 L4 L5 Avg.
Qwen2.5 7B -NL 50.21 34.80 32.75 25.84 14.92 31.71 50.21 52.02 53.04 46.98 42.67 48.98
Granite 8B Code -NL 44.92 36.70 33.13 21.01 13.61 29.88 44.92 50.35 50.36 40.70 39.44 45.15
Fine-grained Instructions: FollowBench (Jiang et al., 2024b) includes instruction tasks where fine-grained constraints are incrementally introduced at each ‘level’. The dataset consists of $55$ levels of constraints applied to different task types - ‘style’, ‘situations’, ‘content’, ‘format’ and ‘examples’. Models are scored by assessing how well all constraints are satisfied for each instruction (hard-constraint satisfaction rate) and by assessing whether a particular constraint is satisfied, regardless of whether co-occurring constraints are satisfied are not (soft-constraint satisfaction rate).
We report the results of this experiment in Table 6. We find that with the exception of Granite 8B Code and Granite 3.1 8B models,***We inspected the results from the Granite models but could not find a pattern that explains this behavior. all other models benefit from training with pseudo-code instructions. Further, performance gains when pseudo-code training is used are especially significant at higher (more complex) levels of constraints.
3.3.2 Code Tasks
Though most modern LLMs are pre-trained on a substantial amount of code data, our training mixture included only a small proportion of code-related tasks (15151515% of the training data). Further, these tasks were treated no differently than other NL tasks. For example, for a code-synthesis task, before attempting the code-task models first need to generate a pseudo-code of the instruction and then create the code requested. Also, given that our SFT training uses a python-like representation for pseudo-code, we investigate how the performance is affected by the inclusion of such pseudo-code using the HumanEvalPack code benchmark  (Muennighoff et al., 2024).
The HumanEvalPack benchmark consists of $33$ main categories of tasks including – Code Repair, Code Explanation, Code Synthesis across $66$ programming languages – Python, JavaScript, Java, Go, C++ and Rust. Following prior work  (Chen et al., 2021; Muennighoff et al., 2024), we evaluate the performance of all code tasks using the Pass@1 metric, which measures the rate at which the first generated solution passes all test cases upon execution.
Table 7: Model performance on low-score HumanEval benchmarks. Metrics: FixD = FixDocs, FixT = FixTests, Synth = Synthesize, ExplSyn = ExplainSynthesize, Py = Python, Oth = Other.
Model FixD-Py FixD-Oth FixT-Py FixT-Oth Synth-Py Synth-Oth ExplSyn-Py ExplSyn-Oth Avg
Llama 3.1 8B -NL 0.439 0.111 0.22 0.066 0.402 0.104 0.268 0.087 0.212
Llama 3.1 8B -PC 0.189 0.039 0.122 0.024 0.262 0.121 0.299 0.109 0.146
Mistral 7B v0.3 -NL 0.177 0.043 0.079 0.029 0.250 0.071 0.220 0.051 0.115
Mistral 7B v0.3 -PC 0.085 0.009 0.030 0.002 0.195 0.080 0.122 0.040 0.071
Qwen2.5 7B -NL 0.720 0.205 0.500 0.160 0.604 0.184 0.543 0.146 0.383
Qwen2.5 7B -PC 0.268 0.154 0.457 0.12 0.055 0.184 0.226 0.095 0.195
Granite 8B Code -NL 0.634 0.200 0.396 0.129 0.537 0.155 0.427 0.133 0.326
Granite 8B Code -PC 0.512 0.145 0.299 0.111 0.372 0.156 0.360 0.132 0.261
Granite 3.1 8B -NL 0.488 0.139 0.256 0.094 0.476 0.144 0.311 0.100 0.251
Granite 3.1 8B -PC 0.110 0.032 0.030 0.013 0.390 0.129 0.140 0.052 0.112
Based on Table 7, we find that pseudo-code (PC) instruction-training using our mixture consistently reduces performance across all models and tasks compared to natural language (NL) tuning. The drop is especially sharp for Python tasks; supporting the hypothesis that our Python-like pseudo-code affects real Python code understanding. Code synthesis is the most affected task, with models like Qwen2.5 7B dropping from 0.604 to 0.055 (Pass@1). Code repair and explanation also degrade, though less severely. This suggests that training with pseudo-code instructions requires careful consideration of the nature and amount of code tasks provided during SFT.
4 Related Work
Our work is most closely related to the creation of intermediate thinking/reasoning representations LLMs. Most often, these are applied at inference time - styles include simple prompts to elicit thinking  (Kojima et al., 2022), instructions that guide LLMs to breakdown tasks and plan  (Xu et al., 2023), structured representation and meta chain-of-thought (eg: Tree Of Thoughts  (Yao et al., 2023a), Graph of Thoughts  (Besta et al., 2024), Table of Thoughts (Sun et al., 2025), Buffer of Thoughts (Yang et al., 2024b)), specialized representations for tasks (eg: Tabular Reasoning  (Wang et al., 2024), Numerical Reasoning  (Chen et al., 2023), Code Generation  (Li et al., 2025), Conditional QA  (Lin et al., 2024)), etc. Additionally, prompting strategies that use code-like representation to aid thinking have also been developed  (Mishra et al., 2023; Li et al., 2024a; Puerto et al., 2024). Our work extends this body of work on the use of code-like representations - specifically, instead of relying on tedious few-shot demonstrations to elicit code-generation during inference, we fine-tune models to do this as part of a ‘thinking’ step before returning the answer. We mirror the approach taken by  Mishra et al. (2023) and  Puerto et al. (2024) of using pseudo-code as ‘instructions’ during inference and do not rely on the use of a code-emulator for line-by-line attempted execution used in Chain-of-Code (Li et al., 2024a). Thus, our models do not require external code-execution environments or emulators can be used as drop-in replacements for their corresponding NL-instruction-tuned model counterparts.
5 Conclusion and Future Work
In this paper experimented with an instruction-tuning strategy that trains models to first re-express the instruction in pseudo-code before attempting to answer the task. We modified a portion of the publicly available TuluV2 dataset  (Ivison et al., 2023) and trained $55$ publicly available base models. We conducted rigorous experiments on a variety of instruction-following tasks and demonstrated the effectiveness of fine-tuning with pseudo-code as compared to only using natural language instructions. Further, we demonstrated that the performance on mathematical as well as common-sense reasoning is, on average, comparable or better than when models are trained only using NL instructions.
Our work opens up interesting directions of future work. Since we wanted to study the effect of using pseudo-code we made sure they do not contain any chain-of-thought or contain other forms of instruction enrichment  (Qi et al., 2024; Lou et al., 2024). Further, our results from Table 2 suggests that for some model-task combinations involving common-sense and reasoning, training with vanilla NL instructions may be better. This suggests it may be useful to explore techniques to combine capabilities such as using mixture-of-experts architectures (Jiang et al., 2024a), model merging techniques  (Yang et al., 2024a), joint NL and NL-PC training mixtures, experimentation with pretraining and more. Further, our instruction-mixture did not use a lot of code-tasks; experimenting with alternative strategies of mixing code-tasks such as including code-tasks without pseudo-code generation and increasing the proportion of such tasks in the mixture could help mitigate some of the issues related to code-task deterioration. Lastly, we believe training models to ‘think’ and ‘plan’ in ‘code’ could enable even more powerful applications where models could exploit frameworks such as LMQL (Beurer-Kellner et al., 2023) for control-flow to support mixed LLM-executable and code-executable programs.
Appendix A Overview of Appendix
Our appendix is organized as follows. We begin by including PC and NL training samples in Appendix Section B. Then we include sample outputs from trained models in Appendix Section C. We then include the quantitative and qualitative results of running the Repair stage in Appendix Section D. We include data specific instances pseudo-code programs used in the creation of training data in Appendix Section E and then provide details on how model outputs are processed for evaluation in Appendix Section F. Finally, we include some details on statistical significance (Appendix Section G) and conclude the Appendix by discussing limitations of our work in Appendix Section H.
Appendix B Training Sample
@@
Appendix C Qualitative Samples
C.1 IFEval
C.1.1 Example 1
@@ C.8.2 Example 2
Appendix D Repair Stage
@@ Appendix E Prompts
Appendix F Model Output Post Processing

### v2

@@ Appendix B Performance on Instructions with Distractors
Appendix C Performance on Code Tasks
Table 8: Model Performance Scores Across All Programming Languages (NL and PC-COT) [^8] The results reported in an earlier version of this paper were affected by a bug in the evaluation code. The issue has been identified and corrected, and all results have been recomputed accordingly.
FixTests FixDocs Synthesize ExplainSynth
Model JS Py C++ Rust JS Py C++ Rust JS Py C++ Rust JS Py C++ Rust Avg
Llama 3.1 8B-NL 0.213 0.220 0.116 0.140 0.390 0.445 0.262 0.165 0.341 0.372 0.268 0.189 0.049 0.067 0.037 0.049 0.208
Llama 3.1 8B-PC 0.104 0.226 0.146 0.006 0.317 0.427 0.280 0.159 0.378 0.445 0.299 0.213 0.079 0.049 0.037 0.037 0.200
Mistral 7B v0.3-NL 0.012 0.085 0.067 0.110 0.043 0.159 0.055 0.226 0.134 0.232 0.213 0.085 0.055 0.043 0.061 0.018 0.097
Mistral 7B v0.3-PC 0.043 0.079 0.049 0.043 0.189 0.165 0.146 0.067 0.213 0.250 0.195 0.067 0.067 0.073 0.067 0.006 0.118
Qwen2.5 7B-NL 0.476 0.500 0.329 0.146 0.622 0.683 0.616 0.183 0.543 0.610 0.488 0.183 0.159 0.140 0.085 0.183 0.372
Qwen2.5 7B-PC 0.134 0.500 0.335 0.195 0.677 0.683 0.573 0.476 0.628 0.652 0.500 0.494 0.183 0.171 0.073 0.146 0.401
Granite 8B Code-NL 0.232 0.378 0.274 0.226 0.518 0.622 0.524 0.201 0.476 0.476 0.415 0.159 0.079 0.098 0.079 0.055 0.301
Granite 8B Code-PC 0.323 0.409 0.274 0.262 0.616 0.659 0.573 0.463 0.561 0.555 0.427 0.402 0.091 0.073 0.085 0.061 0.365
Granite 3.1 8B-NL 0.268 0.256 0.195 0.226 0.372 0.470 0.390 0.250 0.415 0.463 0.390 0.122 0.079 0.067 0.024 0.030 0.251
Granite 3.1 8B-PC 0.195 0.293 0.189 0.165 0.390 0.457 0.372 0.244 0.366 0.415 0.390 0.207 0.030 0.043 0.030 0.030 0.239
Since our SFT training includes Python-like pseudo-code, we investigate its impact on HumanEvalPack code benchmarks (Muennighoff et al., 2024). HumanEvalPack consists of three main categories—Code Repair, Code Explanation, and Code Synthesis—across six programming languages: Python, JavaScript, Java, Go, C++, and Rust. Although our training mixture included only a small proportion of code-related tasks (15% of the training data), most modern LLMs are pre-trained on substantial amounts of code.
We evaluate models on four HumanEvalPack tasks that test code-generation reasoning and robustness. FixTests requires repairing incorrect code using only failing unit tests. FixDocs removes test supervision, challenging the model to fix code from natural language documentation. Synthesize is the standard code-generation task, producing programs from problem descriptions. ExplainSynthesize evaluates whether models can generate correct implementations from intermediate explanations, testing abstract reasoning.
Table 8 shows that, on average, training with pseudo-code instructions improves performance over natural-language prompting across HumanEvalPack tasks. Granite 8B Code exhibits the largest average gain, increasing from 0.301 (NL) to 0.365 (pseudo-code), with especially large improvements on FixDocs and Synthesize in Rust. While gains vary by model and task, these results suggest that re-expressing instructions in a pseudo-code form can positively affect code-generation performance, even when the representation is limited to instruction specification rather than solution guidance.
Appendix D Training Sample
@@
Appendix E Qualitative Samples
E.1 IFEval
E.1.1 Example 1
@@ E.8.2 Example 2
Appendix F Repair Stage
@@ Appendix G Prompts
Figure 20: Pseudocode for Code Task
def generate_code_grounded_on_instruction(instruction_with_input: str) -> str:
    this function generates the code snippet/function for the given instruction and input
    Parameters:
        str: instruction_with_input
    Returns:
        str: code
    # this function extracts and returns the task instruction from the given input string
    task = extract_task_info_from_instruction(instruction_with_input)
    # this function extracts and returns the input parameters if present in the instruction, else returns None
    input = extract_input_from_instruction(instruction_with_input)
    # this function extracts and returns the programming language in which the code snippet needs to be written.
    # If no programming language information is present, then returns pseudocode
    programming_language = extract_programming_language_from_instruction(instruction_with_input)
    code = generate_code(task=task, input=input, programming_language=programming_language)
    return code
>>> generate_code_grounded_on_instruction(Write a Python function ‘generate_integers(a, b)‘ to solve the following problem:
Given two positive integers a and b, return the even digits between a
and b, in ascending order.
For example:
generate_integers(2, 8) => [2, 4, 6, 8]
generate_integers(8, 2) => [2, 4, 6, 8]
generate_integers(10, 14) => []
Figure 21: One shot Prompt for CoT ARC Dataset
Question: Which factor will most likely cause a person to develop a fever?
Options:
A. a leg muscle relaxing after exercise
B. a bacterial population in the bloodstream
C. several viral particles on the skin
D. carbohydrates being digested in the stomach
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a multiple-choice question with 4 options, return the most appropriate option.
    Parameters:
        question_with_options (str): Full question string including options A-D.
    Returns:
        str: The correct option label (’A’, ’B’, ’C’, or ’D’).
    # Step 1: Parse the question and options
    question, options = parse_question_and_options(question_with_options)
    # Step 2: Evaluate each option’s relevance to the question
    scores = {}
    for label, option in options.items():
        scores[label] = evaluate_option_relevance(question, option)
    # Step 3: Select the option with the highest score
    best_option = select_best_option(scores)
    return best_option
[…]
[/PSEUDOCODE]
Answer: B
Question: ...
Figure 22: One shot Prompt for CoT GSM8K and GSM8K Platinum Dataset
Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
[PSEUDOCODE]
def answer_question(question: str) -> int:
    Given a question string, extract relevant numbers and compute the total clips sold.
    Parameters:
        question (str): A natural language math question.
    Returns:
        int: The final computed answer.
    # Step 1: Extract April sales
    april_clips = extract_april_sales(question)
    # Step 2: Extract May multiplier
    may_factor = extract_may_factor(question)
    # Step 3: Calculate May sales
    may_clips = april_clips * may_factor
    # Step 4: Total sales
    total_clips = april_clips + may_clips
    return int(total_clips)
>>> answer_question("Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in
April and May?")
[/PSEUDOCODE]
Answer: 72
Question: ...
Figure 23: One shot Prompt for CoT Hellaswag Dataset
Given a partial sentence and four options 0, 1, 2, and 3. Generate the option number corresponding to the correct ending. Generate pseudocode chain-of-thought
reasoning first and then generate the solution.
Sentence: Getting a haircut: The man in the center is demonstrating a hairstyle on the person wearing the blue shirt. The man in the blue shirt
Options:
0. is standing on the sponge cutting the hair of the person wearing the blue shirt.
1. is doing the hairstyle with his hand and the hairspray.
2. sits on the chair next to the sink.
3. is being shown eye to eye.
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a partial sentence with 4 options (0-3), return the option number that correctly completes the sentence.
    Parameters:
        question_with_options (str): Full question string including sentence and options.
    Returns:
        str: The correct option label (’0’, ’1’, ’2’, or ’3’).
    # Step 1: Parse the sentence and options
    sentence, options = parse_sentence_and_options(question_with_options)
    # Step 2: Evaluate each option for logical and grammatical fit
    scores = {}
    for label, option in options.items():
        scores[label] = evaluate_sentence_completion(sentence, option)
    # Step 3: Select the option with the highest score
    best_option = select_best_option(scores)
    return best_option
>>> answer_question("Getting a haircut: The man in the center is demonstrating a hairstyle on the person wearing the blue shirt. The man in the blue shirt
[…]
[/PSEUDOCODE]
Answer: 2
Sentence: ...
Figure 24: One shot Prompt for CoT PiQA Dataset
Question: how do you indent something?
Options:
0. leave a space before starting the writing
1. press the spacebar
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a question with two options (0 and 1),
    return the correct option as a string ("0" or "1").
    # Step 1: Extract the question text
    question_text = extract_question_text(question_with_options)
    # Step 2: Extract the options
    options = extract_options(question_with_options)
    # Step 3: Use reasoning to determine the correct answer
    correct_answer = infer_correct_option(question_text, options)
    # Step 4: Return only "0" or "1"
    return correct_answer
[…]
[/PSEUDOCODE]
Answer: 0
Question: ...
Figure 25: One shot Prompt for CoT MathQA Dataset
Question: sophia finished 2 / 3 of a book . she calculated that she finished 90 more pages than she has yet to read . how long is her book ?
Options:
a. 229
b. 270
c. 877
d. 266
e. 281
[PSEUDOCODE]
def answer_question(question_with_options: str) -> int:
    Given a math word problem with options, solve it step by step.
    # Step 1: Extract finished fraction from the question
    finished_fraction = extract_finished_fraction(question_with_options)
    # Step 2: Extract difference in pages
    difference_pages = extract_difference_pages(question_with_options)
    # Step 3: Compute remaining fraction
    remaining_fraction = 1 - finished_fraction
    # Step 4: Solve equation for total pages
    total_pages = difference_pages / (finished_fraction - remaining_fraction)
    # Step 5: Extract options from the question
    options = extract_options(question_with_options)
    # Step 6: Match the computed answer with options
    for option in options:
        if option == int(total_pages):
            return option
>>> answer_question("sophia finished 2 / 3 of a book . she calculated that she finished 90 more pages than she has yet to read . how long is her book ?
[…]
[/PSEUDOCODE]
Answer: 270
Question: ...
Figure 26: One shot Prompt for CoT Winogrande Dataset
Given a sentence with one of the word replaced by an underscore(_) and two options 1 and 2. Choose either 1 or 2 which best replaces the underscore. Generate
pseudocode chain-of-thought reasoning first and then generate the solution.
Sentence: He never comes to my home, but I always go to his house because the _ is smaller.
Options:
1. home
2. house
[PSEUDOCODE]
def answer_question(question_with_options: str) -> str:
    Given a sentence with a blank (_) and two options (1 and 2),
    return the correct option as a string ("1" or "2").
    # Step 1: Extract the sentence
    sentence = extract_sentence(question_with_options)
    # Step 2: Extract the options
    options = extract_options(question_with_options)
    # Step 3: Infer the correct replacement using context
    correct_option = infer_correct_replacement(sentence, options)
    # Step 4: Return only "1" or "2"
    return correct_option
[…]
[/PSEUDOCODE]
Answer: 1
Sentence: ...
Figure 27: One shot Prompt for CoT Complex Bench and Follow Bench Dataset
Question: Write a tweet-length description (max 280 characters) of the benefits of learning to code. Avoid all technical jargon, use plain language that anyone
can understand, do not use emojis, and include a real-life example of how coding helps in everyday life.
[PSEUDOCODE]
def answer_question(question: str) -> str:
    Given a prompt asking for a formatted or constrained response,
    extract the constraints and generate a valid answer accordingly.
    Parameters:
        question (str): Instructional prompt.
    Returns:
        str: A response satisfying all the constraints.
    # Step 1: Parse the constraints
    constraints = extract_constraints(question)
    # Step 2: Identify the core task (e.g., benefits of learning to code)
    topic = extract_topic(question)
    # Step 3: Generate a candidate answer for the topic
    candidate = generate_plain_language_response(topic)
    # Step 4: Ensure the candidate satisfies all constraints
    validated_answer = enforce_constraints(candidate, constraints)
    return validated_answer
>>> answer_question("Write a tweet-length description (max 280 characters) of the benefits of learning to code. Avoid all technical jargon, use plain language
that anyone can understand, do not use emojis, and include a real-life example of how coding helps in everyday life.")
[/PSEUDOCODE]
Answer: Learning to code can help you automate boring tasks—like renaming hundreds of files at once. It also builds logical thinking and opens doors to tech jobs.
Anyone can start small, practice daily, and create things that make life easier.
Question: ...
Figure 28: One shot Prompt for CoT IFEval Dataset
Question: List two animals that live in the ocean. Separate your answers with 5 dollar symbols: \$\$\$\$\$
[PSEUDOCODE]
def answer_question(question: str) -> str:
    Given a natural language question, extract constraints (like number of items, type of items,
    and required formatting) and generate a valid answer.
    Parameters:
        question (str): The input question.
    Returns:
        str: The generated answer satisfying both content and formatting constraints.
    # Step 1: Parse what type of items are requested (e.g., animals, colors, cities)
    item_type = extract_item_type(question)  # e.g., "animals that live in the ocean"
    # Step 2: Parse how many items are requested
    num_items = extract_quantity(question)  # e.g., 2
    # Step 3: Extract the required delimiter or format
    delimiter = extract_delimiter(question)  # e.g., "\$\$\$\$\$"
    # Step 4: Retrieve or generate items of the correct type
    items = generate_items(item_type, num_items)  # e.g., ["dolphin", "octopus"]
    # Step 5: Join the items using the specified delimiter
    formatted_answer = delimiter.join(items)
    return formatted_answer
>>> answer_question("List two animals that live in the ocean. Separate your answers with 5 dollar symbols: \$\$\$\$\$")
[/PSEUDOCODE]
Answer: dolphin \$\$\$\$\$ octopus
Question: ...
Figure 29: One shot Prompt for CoT AgentIF Dataset (part 1)
Task : You are an experienced mental health professional speaking directly to the user. Your task is to:
1. Create a safe space by acknowledging their courage in seeking support
2. Analyze their emotional state with clinical precision and genuine empathy
3. Ask targeted follow-up questions to understand their full situation
4. Identify patterns in their thoughts, behaviors, and relationships
5. Assess risk levels with validated screening approaches
6. Help them understand their current mental health in accessible language
7. Validate their experiences without minimizing or catastrophizing
You should address both “analyze emotional state” and “identify patterns in thoughts, behaviors, and relationships” in separate sentences or paragraphs,
using non-overlapping language and observations to avoid redundancy. Always use "you" and "your" when addressing the user. Blend clinical expertise with genuine
warmth and never rush to conclusions.
Please first provide a 2-3 sentence summary of your ideas on the assessment based on the context provided.
Your task
Your task is write the assessment part of the report. Do not include any other parts. Do not use XML tags.
Start your response with: ’## ASSESSMENT Design’.
Below are some context for you to refer to:
Emotional State: I feel constantly on edge and emotionally drained, like I can’t fully relax even when nothing is happening. I notice frequent worry and a
lingering sense that I’m falling behind in life.
Sleep: Around 5–6 hours per night; difficulty falling asleep due to racing thoughts, occasional early-morning awakening.
Stress Level: 8/10
Support System: [’My sister Ananya, who I speak with every few days’, ’A college friend, Rohit, who messages occasionally’, ’A therapist I saw briefly last year
but stopped due to time constraints’]
Recent Changes: I recently took on additional responsibilities at home due to a family member’s illness, while also struggling to meet increasing expectations at
work.
Current Symptoms: [’Muscle tension’, ’Mental fatigue’, ’Reduced motivation’, ’Difficulty making decisions’]
[PSEUDOCODE]
def mental_health_assessment(context_text: str) -> str:
    Generate the ASSESSMENT section of a mental health report from unstructured text input.
    The function parses narrative context and produces a structured clinical assessment.
    Parameters:
        context_text (str): Free-form text describing emotions, stressors, symptoms,
                            sleep, support system, and recent life changes.
    Returns:
        dict: A structured assessment containing summary, emotional analysis,
              pattern identification, risk assessment, psychoeducation, and validation.
    assessment = {}
    emotional_state = extract_emotional_state(context_text)
    stress_level = extract_stress_level(context_text)
    sleep_profile = extract_sleep_information(context_text)
    recent_changes = extract_recent_life_events(context_text)
    symptoms = extract_current_symptoms(context_text)
    support_system = extract_support_system(context_text)
    assessment["summary"] = summarize_overall_presentation(
        emotional_state=emotional_state,
        stress_level=stress_level,
        recent_changes=recent_changes,
        symptoms=symptoms
    assessment["emotional_state_analysis"] = analyze_emotional_state(
        emotions=emotional_state,
        sleep=sleep_profile,
        somatic_and_cognitive_symptoms=symptoms
    assessment["patterns_analysis"] = identify_patterns(
        life_events=recent_changes,
        sleep=sleep_profile,
        supports=support_system,
        symptoms=symptoms
    assessment["risk_assessment"] = assess_risk_level(
        anxiety_intensity=stress_level,
        sleep_disruption=sleep_profile,
        functional_impairment=symptoms
    assessment["psychoeducation"] = explain_mental_state_accessibly(
        emotional_state=emotional_state,
        stressors=recent_changes
Figure 30: One shot Prompt for CoT AgentIF Dataset(part 2)
    assessment["validation"] = validate_user_experience(
        pressures=recent_changes,
        functional_impact=symptoms
    return assessment
>>> mental_health_assessment(context_text)
[/PSEUDOCODE]
Answer: ## ASSESSMENT Design
You are presenting with sustained psychological strain marked by chronic worry, emotional exhaustion, and difficulty disengaging mentally, even during periods
of rest. Your reduced sleep quality, muscle tension, and decision-making fatigue indicate that your stress response has remained activated over an extended
period, rather than resolving naturally.
When examining patterns in your thoughts, behaviors, and relationships, there is a recurring theme of increased responsibility combined with high self-
expectations and limited emotional offloading. You appear to prioritize meeting external demands while minimizing your own needs, which may contribute to
internal pressure, mental fatigue, and reduced motivation. Although you have access to supportive relationships, your engagement with them seems intermittent,
suggesting you may hesitate to rely fully on others during periods of stress.
From a risk assessment standpoint, your stress level (8/10), ongoing sleep disruption, and cognitive fatigue align with moderate anxiety symptoms as reflected
in screening measures such as the GAD-7. There is no indication of acute safety concerns or suicidal ideation at this time; however, the persistence of symptoms
and growing functional impact suggest an elevated vulnerability to burnout or depressive features if current stressors remain unaddressed. Continued monitoring
and re-engagement with structured support would be clinically advisable.
Overall, your current mental health can be understood as the cumulative effect of prolonged role strain and emotional overextension rather than a personal
inadequacy. Your reactions are understandable given the pressures you are managing, and they signal a need for support, recalibration, and restoration rather
than endurance alone.
Task : ...
Appendix H Model Output Post Processing